### v1

# General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

Abstract.
In this work, we explore the application of the Virtual Element Methods for Neumann boundary Optimal Control Problems in saddle point formulation. The method is proposed for arbitrarily polynomial order of accuracy and general polygonal meshes. Our contribution includes a rigorous a priori error estimate that holds for general polynomial degree. On the numerical side, we present (i) an initial convergence test that reflects our theoretical findings, and (ii) a second test case based on a more application-oriented experiment. For the latter test, we focus on the role of VEM stabilization, conducting a detailed experimental analysis, and proposing an alternative structure-preserving strategy to circumvent issues related to the choice of the stabilization parameter.
1. Introduction
In this paper, we consider Virtual Element Methods (VEM) approximation for optimal Neumann boundary control. Optimal Control Problems (OCPs) constrained to partial differential equations (PDEs) are widespread in many practical applications in scientific and engineering contexts, see e.g. [2, 30, 32, 33] and the references therein. The main goal of OCPs is to reach a beneficial behavior of the PDE solution by employing an external variable, the control. The role of the control consists in making the state solution similar to a desired configuration using constrained minimization of a functional. The control is distributed when its action is defined all over the domain, while we talk about boundary control if the control affects the boundary conditions. Due to its coupled nature, the numerical analysis for OCPs constrained to PDEs is a fascinating mathematical challenge. Several works have been proposed with Finite Element Methods, Discontinuous Galerkin Methods, Finite Volumes and hybrid methods: we refer the reader to this far from exhaustive list and the reference therein [18, 21, 24, 27, 42, 43, 44]
Lately, great attention has been paid to VEM, a family of polygonal methods for simulating PDEs. VEM can be applied on arbitrarily polygonal meshes, guaranteeing huge flexibility in terms of geometrical complexity and they might play a crucial role in many real-life scenarios. Indeed, after being conceived in the seminal paper [3], they have been applied to various contexts of interest [5, 7, 13, 19, 20, 23, 41]. The investigation of boundary OCPs through VEM is still limited yet of utmost importance. Indeed, the versatility of arbitrary polygonal meshes, combined with the action of control strategies on complex geometries, might be not only a challenging theoretical topic but also relevant from a practical viewpoint. To our knowledge, $C^{1}$-VEM is presented in [16] for distributed OCPs constrained to elliptic equations and with a constrained state variable. The VEM discretization for distributed control is the topic of several investigations, see e.g. [34, 36, 37, 38, 40]. Some VEM works related to boundary OCPs are [39] for the Dirichlet case with control constraints and [25] for the Neumann case with control constraints and approximation of the first order.
In this contribution, we focus on the unconstrained Neumann boundary control problem in a saddle point structure. This naturally derives from the optimality conditions and considers the coupled nature of the system leading to more robust and stable systems solving for all the variables at once. This formulation is generally more effective than iterative algorithms. To the best of our knowledge, we differentiate from previous contributions in the following terms:
  • the presented method has arbitrary polynomial accuracy;
  • we analyze the problem in saddle point formulation as presented in [28];
  • we prove well-posedness and a priori error estimates for a general polynomial degree and variable coefficients;
  • numerical tests cover high-order schemes;
  • we numerically investigate how the stabilization parameter affects the method’s accuracy.
Moreover, we propose a structure-preserving strategy to circumvent issues related to VEM stabilizations, based on recent developments in VEM theory [8, 9, 10, 12].
The paper is structured as follows: in Section 2 we introduce the linear-quadratic Neumann boundary OCP at the continuous level in a saddle point framework. Section 3 focuses on the VEM approximation of such optimization problem and proves the well-posedness of the saddle point structure. In Section 4 we derive optimal a priori error estimates. Section 5 presents two test cases: a convergence analysis test and a more application-oriented test for which we discuss the role of the stabilization parameter, proposing a workaround strategy using structure-preserving VEM. Finally, conclusions follow in Section 6.
Throughout the work, $(\cdot,\cdot)_{\omega}$ denotes the standard $\mathrm{L}^{2}$ scalar product defined on a generic $\omega\subset\mathbb{R}^{2}$ and inside the proofs, the symbol $C$ denotes any constant independent of the mesh size.
2. Problem Formulation
Let us consider $\Omega\subset\mathbb{R}^{2}$ as an open and bounded regular domain. We denote as $\Gamma_{D}$ and $\Gamma_{N}$ the portions of the boundary $\partial\Omega$ where Dirichlet and Neumann conditions apply. Moreover, we define a control boundary $\Gamma_{C}$, where the control acts in form of Neumann condition. We remark that $\Gamma_{D}\cup\Gamma_{N}\cup\Gamma_{C}=\partial\Omega$ and $\Gamma_{D}\cap\Gamma_{N}\cap\Gamma_{C}=\emptyset$. Moreover, for simplicity we assume $\Gamma_{D}\neq\emptyset$.
The state variable $y$ and the control variable $u$ are considered, respectively, in
$Y=\{y\in H^{1}(\Omega)\;|\;y=0\text{ on }\Gamma_{D}\}\text{ and }U=L^{2}(\Gamma_{C}).$
The case of non-homogeneous boundary conditions can be tackled by standard lifting procedures [31]. Given a desired state $y_{\mathrm{d}}\in Y_{\text{d}}=L^{2}(\Omega_{\text{obs}})$, with $\Omega_{\text{obs}}\subset\Omega$, a forcing term $f\in L^{2}(\Omega)$ and a penalization parameter $\alpha>0$ on the control action, we solve the following problem.
Problem 1.
Find $(y,u)\in Y\times U$ such that
$J(y,u)=\min_{(w,v)\in Y\times U}J(w,v)=\min_{(w,v)\in Y\times U}\frac{1}{2}\left\lVert{w-y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(\Omega_{\text{obs}})}^{2}+\frac{\alpha}{2}\left\lVert{v}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}\,,$
subject to the following PDE
(1) $\begin{cases}-\nabla\cdot(\mathcal{K}\nabla w)+\gamma w=f&\text{in }{\Omega},\\ \displaystyle w=0&\text{on }{\Gamma_{D}},\\ \displaystyle\frac{\partial w}{\partial n}=v&\text{on }{\Gamma_{C}},\vspace{0.5mm}\\ \displaystyle\frac{\partial w}{\partial n}=0&\text{on }{\Gamma_{N}},\\ \end{cases}$
where $\mathcal{K}$ and $\gamma$ are a uniformly elliptic tensor valued function and a scalar function, respectively, both considered sufficiently smooth.
Problem 1 can be expressed in weak formulation as find the pair $(y,u)\in Y\times U$ which minimizes $J(y,u)$ and satisfies
(2) ${a(y,q)}=\left(u,q\right)_{\Gamma_{C}}+\left(f,q\right)_{\Omega}\quad\forall q\in Y,$
where $a:Y\times Y\rightarrow\mathbb{R}$ is given by
(3) $a(y,q)\vcentcolon=\left(\mathcal{K}\nabla y,\nabla q\right)_{\Omega}+\left(\gamma y,q\right)_{\Omega}\,\quad\forall y,q\in Y.$
Since $a(\cdot,\cdot)$ is continuous and coercive and $\left(u,q\right)_{\Gamma_{C}}$ is continuous, (2) is well-posed [31]. Let $\left\lvert\left\lvert\left\lvert\cdot\right\rvert\right\rvert\right\rvert_{Y}$ denote the norm induced by ${a}$, i.e. for each $w\in Y$
(4) $\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}^{2}{\;\vcentcolon=\;}\left\lVert{\sqrt{\mathcal{K}}\nabla w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}+\left\lVert{\sqrt{\gamma}w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}\,.$
Remark 1.
For any $w\in Y$, there exists $C>0$ independent of $h$ such that
(5) $\left\lVert{w}\right\rVert_{\mathrm{H}^{1}\!(Y)}\leq C\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}$
holds true, due to a Poincarè inequality.
We define the weak minimization problem as
[…]
Considering $\alpha>0$ and thanks to the assumptions over the bilinear forms $a(\cdot,\cdot)$ and $\left(u,q\right)_{\Gamma_{C}}$, the minimization problem is well-posed and a Lagrangian approach can be exploited to solve it [22]. First, we define an adjoint variable $p\in Y$ and the Lagrangian functional as
(7) $\mathscr{L}(y,u,p)=J(y,u)+\displaystyle a(y,p)-\left(u,p\right)_{\Gamma_{C}}-\left(f,p\right)_{\Omega}.$
The pair $(y,u)$ minimizing (6), is obtained solving the optimality system that finds $(y,u,p)\in Y\times U\times Y$ such that
(8) $\begin{cases}D_{y}\mathscr{L}(y,u,p)[z]=0&\forall z\in Y,\\ D_{u}\mathscr{L}(y,u,p)[v]=0&\forall v\in U,\\ D_{p}\mathscr{L}(y,u,p)[q]=0&\forall q\in Y,\\ \end{cases}$
where $D_{y},D_{u}$ and $D_{p}$ denotes the differentiation with respect to the three variables, respectively. In this way, the constrained optimization problem is reframed in an unconstrained optimization and the minimum of (6) is represented by the state and control components of the stationary point of (7). The strong formulation of the optimality system reads:
Problem 2.
Find $(y,u,p)\in Y\times U\times Y$ such that
(9) $\begin{cases}\displaystyle y\boldsymbol{\mathbbm{1}}_{\{\Omega_{\text{obs}}\}}-\nabla\cdot(\mathcal{K}\nabla p)+\gamma p=y_{\mathrm{d}}\boldsymbol{\mathbbm{1}}_{\{\Omega_{\text{obs}}\}}&\text{ in }\Omega,\quad\text{{adjoint equation}}\\ \alpha u-p=0&\text{ in }\Gamma_{C},\;\;\text{{optimality condition}}\\ -\nabla\cdot(\mathcal{K}\nabla y)+\gamma y=f&\text{ in }\Omega,\quad\text{{state constraint}}\\ \displaystyle\frac{\partial y}{\partial n}=u\quad\text{and}\quad\frac{\partial p}{\partial n}=0&\text{ on }\Gamma_{C},\vspace{2mm}\\ \displaystyle\frac{\partial y}{\partial n}=\frac{\partial p}{\partial n}=0&\text{ on }\Gamma_{N},\\ y=p=0&\text{ on }\Gamma_{D},\\ \end{cases}$
where $\boldsymbol{\mathbbm{1}}_{\{\Omega_{\text{obs}}\}}$ is the indicator function of the observation domain $\Omega_{\text{obs}}$.
Let us consider $\boldsymbol{x}:=(y,u)\in X$, where $X$ is the product space between $Y$ and $U$ equipped with the scalar product
(10) $\left(\boldsymbol{x}_{1},\boldsymbol{x}_{2}\right)_{X}=\left(y_{1},y_{2}\right)_{\Omega}+\left(u_{1},u_{2}\right)_{\Gamma_{C}}\,,$
for any $\boldsymbol{x}_{1}=(y_{1},u_{1})$ and $\boldsymbol{x}_{2}=(y_{2},u_{2})$ in $X$, and with the norm
(11) $\left\lVert{\boldsymbol{x}}\right\rVert_{X}^{2}=\left\lvert\left\lvert\left\lvert y\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{u}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2},$
for any $\boldsymbol{x}=(y,u)$.
Then, introducing the bilinear forms ${\mathcal{A}}:X\times X\to\mathbb{R}$ and ${\mathcal{B}}:Y\times X\to\mathbb{R}$ given by
(12) ${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}\right)=\left(y,w\right)_{\Omega_{\text{obs}}}+\alpha\left(u,v\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}=(y,u),\boldsymbol{\varsigma}=(w,v)\in X\,,$
(13) ${\mathcal{B}}\left(q,\boldsymbol{x}\right)=a(y,q)-\left(u,q\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}\in X,\,\forall q\in Y\,.$
and the functional $\mathcal{D}\in X^{*}$
[…]
The weak form of Problem 2 has the following saddle point structure:
Problem 3.
Find $(\boldsymbol{x},p)\in X\times Y$ such that
[…]
By classical well-posedness results Problem 3 has a unique solution and, moreover, it is equivalent to Problem 1, see e.g. [11].
3. VEM Numerical Discretization
In this section, we introduce the VEM approximation of Problem 3.
Let $\mathcal{M}_{h}$ denote a polygonal tessellation of $\Omega$ conforming to $\Omega_{\text{obs}}$. $E$ denotes a general polygon of $\mathcal{M}_{h}$ with diameter $h_{E}$ and let $h$ be the maximum diameter over $\mathcal{M}_{h}$. Let $e$ denote an edge of $\mathcal{M}_{h}$, and $\mathcal{I}_{h}$ be the set of edges lying on $\Gamma_{C}$. We assume that $\mathcal{M}_{h}$ fulfills the following assumptions [6, 15]: $\exists\kappa>0$ such that
  1. A.1
    for all $E\in\mathcal{M}_{h}$, $E$ is star-shaped with respect to a ball of radius $\rho\geq\kappa h_{E}$;
  2. A.2
    for all edges $e$ of $\mathcal{M}_{h}$, $\left\lvert{e}\right\rvert\geq\kappa h_{E}$.
For any given $E\in\mathcal{M}_{h}$, let $\mathbb{P}_{k}\!\left(E\right)$ be the space of polynomials of degree up to $k$ defined on $E$. Let $\Pi^{\nabla,E}_{k}:\mathrm{H}^{1}\!(E)\rightarrow\mathbb{P}_{k}\!\left(E\right)$ be the $\mathrm{H}^{1}$-orthogonal projection defined up to a constant by the orthogonality condition:
(16) $\forall\,w\in\mathrm{H}^{1}\!(E),\quad\left(\nabla\left(\Pi^{\nabla,E}_{k}w-w\right),\nabla z_{k}\right)_{E}=0\;\;\forall\,z_{k}\in\mathbb{P}_{k}\!\left(E\right).$
In order to define $\Pi^{\nabla,E}_{k}$ uniquely, we set $\int_{\partial E}\left(\Pi^{\nabla,E}_{k}u-u\right)\,\mathrm{d}s=0$. For any given $E\in\mathcal{M}_{h}$, let $Y_{h}(E)$ be the local Virtual Element Space:
$Y_{h}(E)=\{w_{h}\in\mathrm{H}^{1}\!(E)\colon\Delta w_{h}\in\mathbb{P}_{k}\!\left(E\right),\,w_{h_{|e}}\in\mathbb{P}_{k}\!\left(e\right)\,\forall e\subset\partial E,\,w_{h}\in C^{0}\!\left(\partial E\right),\\ \left(w_{h},z_{k}\right)_{E}=\left(\Pi^{\nabla,E}_{k}w_{h},z_{k}\right)_{E}\;\forall z_{k}\in\mathbb{P}_{k}\!\left(E\right)\setminus\mathbb{P}_{k-2}\!\left(E\right)\}\,.$
We recall that the degrees of freedom of this space are (see [1, 4])
  • The values of $v$ at the vertices of $E$,
  • If $k\geq 2$, the values at $k-1$ internal distinct points on each edge $e\subset\partial E$,
  • If $k\geq 2$, the scaled moments $\frac{1}{\left\lvert{E}\right\rvert}\int_{E}v\,m_{i}$, where $\{m_{i}\}_{i=1}^{\dim\mathbb{P}_{k-2}\!\left(E\right)}$ is a basis of $\mathbb{P}_{k-2}\!\left(E\right)$.
Moreover, we denote the global discrete space as
$Y_{h}\vcentcolon=\{w_{h}\in Y\colon w_{h_{|E}}\in Y_{h}(E),w_{h}=0\text{ on }\Gamma_{D}\}\,,$
and we equipped it with the norm $\left\lvert\left\lvert\left\lvert\cdot\right\rvert\right\rvert\right\rvert_{Y}$, considering also the following notation for the local norm on each $E\in\mathcal{M}_{h}$
(17) $\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}=\left\lVert{\sqrt{\mathcal{K}}\nabla w}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\sqrt{\gamma}w}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,,$
for each $w\in Y(E)$. The local approximation space for $U$ is simply defined as follows
(18) $\leavevmode\nobreak\ U_{h}(E){\;\vcentcolon=\;}\left\{u_{h}\in U:u_{h}\in\mathbb{P}_{1}\!\left(e\right)\;\forall e\in\mathcal{I}_{h}\,,u_{h}\in C^{0}\!\left(\Gamma_{C}\right)\right\},$
and it is equipped with the $\mathrm{L}^{2}$-norm on $\Gamma_{C}$. To define the discrete formulation of Problem 3, let $\Pi^{0,E}_{k-1}$ be the local $\mathrm{L}^{2}$-orthogonal projector into the space of polynomials of degree up to $k-1$, i.e., for each $w\in\mathrm{L}^{2}\!(E)$, $\int_{E}(\Pi^{0,E}_{k-1}w-w)=0$ and, for each $\boldsymbol{w}=\left(w_{1},w_{2}\right)^{\intercal}\in\left[\mathrm{L}^{2}\!\left(E\right)\right]^{2}$, $\Pi^{0,E}_{k-1}\boldsymbol{w}=\left(\Pi^{0,E}_{k-1}w_{1},\Pi^{0,E}_{k-1}w_{2}\right)^{\intercal}$. Notice that $\Pi^{\nabla,E}_{k}$ and $\Pi^{0,E}_{k-1}$ are computable for any $w_{h}\in Y_{h}(E)$ exploiting the degrees of freedom.
For each $E\in\mathcal{M}_{h}$, let ${a}^{E}_{h}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$ and ${m}^{E}_{h}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$ be the local bilinear forms defined as
(19) ${a}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\mathcal{K}\Pi^{0,E}_{k-1}\nabla w_{h},\Pi^{0,E}_{k-1}\nabla q_{h}\right)_{E}+\left\lVert{\mathcal{K}}\right\rVert_{\infty}S^{E}\left(w_{h}-\Pi^{\nabla,E}_{k}w_{h},q_{h}-\Pi^{\nabla,E}_{k}q_{h}\right)\\ +\left(\gamma\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}+\left\lVert{\gamma}\right\rVert_{\infty}\left\lvert{E}\right\rvert S^{E}\left(w_{h}-\Pi^{\nabla,E}_{k}w_{h},q_{h}-\Pi^{\nabla,E}_{k}q_{h}\right)$
and
(20) ${m}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}\,,$
for each $w_{h},q_{h}\in Y_{h}(E)$. The bilinear form $S^{E}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$, called stabilization, is any symmetric positive definite bilinear form to be chosen to verify that $\exists s_{\ast},s^{\ast}>0$ independent of $h_{E}$ such that
(21) $s_{\ast}\left(\nabla w_{h},\nabla w_{h}\right)_{E}\leq S^{E}\left(w_{h},w_{h}\right)\leq s^{\ast}\left(\nabla w_{h},\nabla w_{h}\right)_{E},$
for each $w_{h}\in Y_{h}(E)$ verifying $\Pi^{\nabla,E}_{k}w_{h}=0$.
Remark 2.
A common choice for $S^{E}$ is the dofi-dofi stabilizing bilinear form $S^{E}\colon Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$, defined such that, denoting by $\chi^{E}(v_{h})$ the vector of degrees of freedom of $v_{h}$ on $E$ and by $\sigma$ an arbitrary-chosen positive constant,
(22) $S^{E}\left(u_{h},v_{h}\right)=\sigma\;\chi^{E}(u_{h})\cdot\chi^{E}(v_{h})\quad\forall u_{h},v_{h}\in Y_{h}(E)\,.$
We report the following classical result [4].
Lemma 1.
There exists $\eta_{\ast}\,,\eta^{\ast}\,,\mu^{\ast}>0$, independent of $h_{E}$, such that for each $w_{h}\in Y_{h}(E)$
(23) $\eta_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}\leq{a}^{E}_{h}\left(w_{h},w_{h}\right)\leq\eta^{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}$
and
(24) ${m}^{E}_{h}\left(w_{h},w_{h}\right)\leq\mu^{\ast}\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,.$
We define the global bilinear forms such that for each $w_{h},q_{h}\in Y_{h}$,
${a}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(w_{h},q_{h}\right)\text{ and }{m}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}{m}^{E}_{h}\left(w_{h},p_{h}\right)\,.$
Let us now write the VEM approximation of Problem 3. Let us consider $\boldsymbol{x}_{h}\vcentcolon=(y_{h},u_{h})\in X_{h}$, where $X_{h}\vcentcolon=Y_{h}\times U_{h}\subset X$ equipped with the norm $\left\lVert{\cdot}\right\rVert_{X}$, i.e. $\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}^{2}=\left\lvert\left\lvert\left\lvert y_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$ , and the bilinear forms ${\mathcal{A}}_{h}:X_{h}\times X_{h}\to\mathbb{R}$ and ${\mathcal{B}}_{h}:Y_{h}\times X_{h}\to\mathbb{R}$ given by
${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)={m}_{h}\left(y_{h},w_{h}\right)+\alpha\left(u_{h},v_{h}\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}_{h}=(y_{h},u_{h}),\boldsymbol{\varsigma}_{h}=(w_{h},v_{h})\in X_{h}\,,$
${\mathcal{B}}_{h}\left(q_{h},\boldsymbol{x}_{h}\right)={a}_{h}\left(y_{h},q_{h}\right)-\left(u_{h},q_{h}\right)_{\Gamma_{C}}\;\forall\boldsymbol{x}_{h}\in X_{h},\,\forall q_{h}\in Y_{h}\,.$
and the functional $\mathcal{D}_{h}\in X_{h}^{*}$
$_{h}{}{X_{h}^{*},X_{h}}=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},\Pi^{0,E}_{k}w_{h}\right)_{E}+\left(0,v_{h}\right)_{\Gamma_{C}}\,.$
Hence, we obtain the following discrete saddle point system.
Problem 4.
Find $(\boldsymbol{x}_{h},p_{h})\in X_{h}\times Y_{h}$ such that
[…]
Proposition 1.
The bilinear forms ${\mathcal{A}}_{h}$ and ${\mathcal{B}}_{h}$ are continuous and satisfy the following conditions
  • $\circ$
    ${\mathcal{A}}_{h}$ is symmetric,
  • $\circ$
    ${\mathcal{A}}_{h}$ is coercive on $X_{h,0}\vcentcolon=\{\boldsymbol{\varsigma}\in X_{h}:{\mathcal{B}}_{h}\left(q_{h},\boldsymbol{\varsigma}_{h}\right)=0\;\forall q_{h}\in Y_{h}\}$, i.e. $\exists\alpha_{\ast}>0$ independent of $h$ such that
    (26) ${\mathcal{A}}_{h}\left(\boldsymbol{\varsigma}_{h},\boldsymbol{\varsigma}_{h}\right)\geq\alpha_{\ast}\,\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}^{2}\quad\forall\boldsymbol{\varsigma}_{h}\in X_{h,0}.$
  • $\circ$
    ${\mathcal{B}}_{h}$ satisfies the inf-sup condition, i.e. $\exists\beta_{\ast}>0$ independent of $h$ such that
    (27) $\inf_{p_{h}\in Y_{h}}\sup_{\boldsymbol{\varsigma}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)}{\left\lvert\left\lvert\left\lvert p_{h}\right\rvert\right\rvert\right\rvert_{Y}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}}\geq\beta_{\ast}\,.$
Then, Problem 4 admits a unique solution and is equivalent to the minimization of
(28) $J_{h}(y_{h},u_{h})=\frac{1}{2}\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}y_{h}-y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\frac{\alpha}{2}\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2},$
subject to ${a}_{h}\left(y_{h},q_{h}\right)=\left(u_{h},q_{h}\right)_{\Gamma_{C}}+\sum_{E\in\mathcal{M}_{h}}\left(f,\Pi^{0,E}_{k}q_{h}\right)_{E}$.
Proof.
It is clear that ${\mathcal{A}}_{h}$ is symmetric and semi-definite positive by the definitions of scalar products. Moreover, ${\mathcal{A}}_{h}$ is continuous, since $\boldsymbol{x}_{h}=(y_{h},u_{h})$, $\boldsymbol{\varsigma}_{h}=(q_{h},v_{h})\in X_{h}$
$|{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)|$ $\leq\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}y_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{\Pi^{0,E}_{k}q_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}+\alpha\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
$\leq\left\lVert{y_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\left\lVert{q_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}+\alpha\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
$\leq\left(1+\alpha\right)\left(\left\lVert{y_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}+\left\lVert{u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\right)\left(\left\lVert{q_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}+\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\right)$
$\leq C\left(1+\alpha\right)\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}\,,$
where we exploit Remark 1. We now prove the coercivity of ${\mathcal{A}}_{h}$ on $X_{h,0}$. Given $\boldsymbol{\varsigma}_{h}\in X_{h,0}$ we have
${\mathcal{B}}_{h}\left(q_{h},\boldsymbol{\varsigma}_{h}\right)=0\quad\forall q_{h}\in Y_{h}\quad\text{if and only if}\quad{a}_{h}\left(w_{h},q_{h}\right)=(v_{h},q_{h})_{\Gamma_{C}}\quad\forall q_{h}\in Y_{h}.$
Using Lemma 1, summing over all the elements of $\mathcal{M}_{h}$ and calling $\eta_{\ast}$ the global coercivity constant with abuse of notation, applying the trace inequality $\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq C_{T}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$ we obtain
(29) $\eta_{\ast}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}\leq{a}_{h}\left(w_{h},w_{h}\right)=\left(v_{h},w_{h}\right)_{\Gamma_{C}}$ $\leq\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
$\leq C_{T}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\,,$
hence
(30) $\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\frac{C_{T}}{\eta_{\ast}}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}.$
Thus,
${\mathcal{A}}_{h}\left(\boldsymbol{\varsigma}_{h},\boldsymbol{\varsigma}_{h}\right)$ $=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\alpha\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}+\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}$
$\geq\frac{\alpha}{2}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}+\frac{\alpha\eta_{\ast}^{2}}{2C^{2}_{T}}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}$
$\geq\frac{\alpha}{2}\min\left(1,\frac{\eta_{\ast}^{2}}{C^{2}_{T}}\right)\left(\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2}\right)\geq\frac{\alpha}{2}\min\left(1,\frac{\eta_{\ast}^{2}}{C^{2}_{T}}\right)\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}^{2}.$
Furthermore, ${\mathcal{B}}_{h}$ is continuous on $Y_{h}\times X_{h}$ by definition. Moreover, ${\mathcal{B}}_{h}$ is inf-sup stable, indeed applying Lemma 1 we have:
$\sup_{0\neq\boldsymbol{x}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(q_{h},\boldsymbol{x}_{h}\right)}{\left\lVert{\boldsymbol{x}_{h}}\right\rVert_{X}}$ $=\sup_{0\neq(y_{h},u_{h})\in Y_{h}\times U_{h}}\frac{{a}_{h}\left(y_{h},q_{h}\right)-(u_{h},q_{h})_{\Gamma_{C}}}{\sqrt{\left\lvert\left\lvert\left\lvert{y}_{h}\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{{v}_{h}}\right\rVert_{U_{h}}^{2}}}$
$\geq\frac{{a}_{h}\left(q_{h},q_{h}\right)}{\left\lvert\left\lvert\left\lvert{q}_{h}\right\rvert\right\rvert\right\rvert_{Y}}\geq\eta_{\ast}\left\lvert\left\lvert\left\lvert{q}_{h}\right\rvert\right\rvert\right\rvert_{Y},$
choosing $(y_{h},u_{h})$ as $(q_{h},0)$ to bound the supremum from below. The arbitrariness of $q_{h}$ concludes the proof. ∎
4. Error Estimates
In order to derive optimal a priori error estimates, we have to provide the polynomial approximation properties and the interpolation error estimates, presented in Lemmas 2, 3 and 4, respectively. The results about polynomial approximation can be led back to classical results ([14]), while the VEM interpolation error estimate is proved in [17] and the interpolation on $U_{h}$ derives from standard results of one-dimensional FEM interpolation.
Lemma 2.
Under assumption A.1, given $\Pi^{0,E}_{k}$ be the local $\mathrm{L}^{2}$-orthogonal projection into $\mathbb{P}_{k}\!\left(E\right)$ and $\Pi^{\nabla,E}_{k}$ $\mathrm{H}^{1}$-orthogonal projection defined by (16), then, there exists $C>0$, independent of $h_{E}$, such that $\forall\phi\in\mathrm{H}^{s}\!(E)$
(31) $\left\lvert{\phi-\Pi^{0,E}_{k}\phi}\right\rvert_{m,E}\leq Ch_{E}^{s-m}\left\lvert{\phi}\right\rvert_{s,E},\;\mbox{ with }m\leq s\leq k+1\,,$
and if $s\geq 1$
(32) $\left\lvert{\phi-\Pi^{\nabla,E}_{k}\phi}\right\rvert_{m,E}\leq Ch_{E}^{s-m}\left\lvert{\phi}\right\rvert_{s,E},\;\mbox{ with }m\leq s\leq k+1\,.$
Lemma 3.
Under assumptions A.1 and A.2, let $w\in\mathrm{H}^{s+1}\!(\Omega)$, $0\leq s\leq k$, there exists $C>0$ such that $\forall\,h$, $\exists w_{I}\in Y_{h}$ satisfying
(33) $\left\lVert{w-w_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}+h\left\lVert{\nabla(w-w_{I})}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\leq Ch^{s+1}\left\lvert{w}\right\rvert_{s+1,\Omega},$
where $C$ is a positive constant, independent of $h$.
Lemma 4.
Under assumptions A.1 and A.2, let $v\in\mathrm{H}^{r}\!(\Gamma_{C})$, $10\leq r\leq k+$ 1. Then there exists a function $v_{I}\in U_{h}$ and a constant $C>0$ independent of $h$ such that
(34) $\left\lVert{v-v_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq Ch^{r}\left\lvert{v}\right\rvert_{r,\Gamma_{C}}\,.$
In order to prove Theorem 1, we need this additional auxiliary result.
Lemma 5.
Let $((y,u),p)\in X\times Y$ be the solution to Problem 3 and let $((y_{h},u_{h}),p_{h})\in X_{h}\times Y_{h}$ be the solution to Problem 4. Moreover, let $u_{I}\in U_{h}$ as in Lemma 4. Then there esists $y_{{\mathcal{B}}_{h}}\in Y_{h}$ such that
(35) ${\mathcal{B}}_{h}\left(q_{h},\left(y_{{\mathcal{B}}_{h}}-y_{h},u_{I}-u_{h}\right)\right)=0\quad\forall q_{h}\in Y_{h}\,.$
Moreover, if $y\in\mathrm{H}^{s+1}\!(\Omega)$, $u\in\mathrm{H}^{r}\!(\Gamma_{C})$, $f\in\mathrm{H}^{t}\!(\Omega)$ with $0\leq s\leq k$, $10\leq r\leq k+$ 1 and $10\leq t\leq k+$ 1, then there exists $C>0$ independent of $h$ such that
(36) $\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq C\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\,.$
Proof.
Let $y_{{\mathcal{B}}_{h}}\in Y_{h}$ be the unique solution to
(37) ${a}_{h}\left(y_{{\mathcal{B}}_{h}},q_{h}\right)=\left(u_{I},q_{h}\right)_{\Gamma_{C}}+\sum_{E\in\mathcal{M}_{h}}\left(\Pi^{0,E}_{k}f,q_{h}\right)\quad\forall q_{h}\in Y_{h}\,.$
Notice that, by definition of ${\mathcal{B}}_{h}$ (3), we have
${\mathcal{B}}_{h}\left(q_{h},(y_{{\mathcal{B}}_{h}},u_{I})\right)=\sum_{E\in\mathcal{M}_{h}}\left(\Pi^{0,E}_{k}f,q_{h}\right)={\mathcal{B}}_{h}\left(q_{h},(y_{h},u_{h})\right)\,,$
thus (35) holds true. Moreover, let $y_{I}\in Y_{h}$ be the function satisfying (33), then by a triangle inequality we have
$\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq\left\lvert\left\lvert\left\lvert y-y_{I}\right\rvert\right\rvert\right\rvert_{Y}+\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\leq Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}\,,$
where $C>0$ is independent of $h$. Regarding the second term above, let $e_{h}=y_{I}-y_{{\mathcal{B}}_{h}}$. Exploiting (23), (37) and the second equation of Problem 3 we get
(38) $\begin{split}\left\lvert\left\lvert\left\lvert y_{I}-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}^{2}&\leq C{a}_{h}\left(y_{I}-y_{{\mathcal{B}}_{h}},e_{h}\right)\\ &=C\Big{(}{a}_{h}\left(y_{I},e_{h}\right)-{a}\left(y,e_{h}\right)\\ &\qquad\qquad+\sum_{E\in\mathcal{M}_{h}}\left(f-\Pi^{0,E}_{k}f,e_{h}\right)_{E}+\left(u-u_{I},e_{h}\right)_{\Gamma_{C}}\Big{)}\,.\end{split}$
The terms above are estimated separately, as follows. By the continuity of ${a}$ and ${a}_{h}$, a triangle inequality and estimates (33) and (32), we get
(39) $\begin{split}{a}_{h}\left(y_{I},e_{h}\right)-{a}\left(y,e_{h}\right)&=\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(y_{I}-\Pi^{0,E}_{k}y,e_{h}\right)+{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)\\ &\qquad-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)+{a}^{E}\left(\Pi^{0,E}_{k}y-y,e_{h}\right)\\ &\leq\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)\\ &\qquad+C\Big{(}\sum_{E\in\mathcal{M}_{h}}\left\lvert\left\lvert\left\lvert y-\Pi^{0,E}_{k}y\right\rvert\right\rvert\right\rvert_{Y(E)}\\ &\qquad\qquad\qquad+\left\lvert\left\lvert\left\lvert y-y_{I}\right\rvert\right\rvert\right\rvert_{Y(E)}\Big{)}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\\ &\leq\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)\\ &\qquad\qquad\qquad\qquad\qquad+Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.\end{split}$
Moreover, the following estimate can be proved by proceeding as in [4, Lemma 5.5]:
$\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(\Pi^{0,E}_{k}y,e_{h}\right)-{a}^{E}\left(\Pi^{0,E}_{k}y,e_{h}\right)\leq Ch^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$
Next, the term of (38) depending on $f$ is estimated exploiting the definition of $\Pi^{0,E}_{k}$, a Cauchy-Schwarz inequality and estimate (31):
$\begin{split}\sum_{E\in\mathcal{M}_{h}}\left(f-\Pi^{0,E}_{k}f,e_{h}\right)_{E}&=\sum_{E\in\mathcal{M}_{h}}\left(f-\Pi^{0,E}_{k}f,e_{h}-\Pi^{0,E}_{k}e_{h}\right)_{E}\\ &\leq\sum_{E\in\mathcal{M}_{h}}\left\lVert{f-\Pi^{0,E}_{k}f}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{e_{h}-\Pi^{0,E}_{k}e_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\\ &\leq Ch^{t}\left\lvert{f}\right\rvert_{t,\Omega}\cdot h\left\lVert{\nabla e_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\leq Ch^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.\end{split}$
Finally, the last term in (38) is bounded by a Cauchy-Schwarz inequality, the estimate (34) and a trace inequality:
$\begin{split}\left(u-u_{I},e_{h}\right)_{\Gamma_{C}}&\leq\left\lVert{u-u_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{e_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\leq Ch^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\left\lVert{e_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\\ &\leq Ch^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\left\lVert{\nabla e_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\\ &\leq Ch^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.\end{split}$
Using the above estimates in (38) and simplifying $\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}$, we get (36).
Theorem 1.
Let $(\boldsymbol{x},p)=((y,u),p)\in\left(\mathrm{H}^{s+1}\!(\Omega)\cap\mathrm{H}^{1}_{0}\!(\Omega)\right)\times\mathrm{H}^{r}\!(\Gamma_{C})\times\left(\mathrm{H}^{s+1}\!(\Omega)\cap\mathrm{H}^{1}_{0}\!(\Omega)\right)$ be the solution to Problem 3, with $0\leq s\leq k$ and $10\leq r\leq k+$ 1. Let $f\in\mathrm{H}^{t}\!(\Omega)$ and $y_{\mathrm{d}}\in\mathrm{H}^{\tau}\!(\Omega_{\text{obs}})$ be the right-hand side of Problem 3, with $10\leq t\leq k+$ 1 and $10\leq\tau\leq k+$ 1. Then, there exists a constant $C$ independent of $h$ such that the unique solution $((y_{h},u_{h}),p_{h})\in X_{h}\times Y_{h}$ to Problem 4 satisfies the following error estimate:
(40) $\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}+\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}&\leq C\big{(}h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\\ &\qquad\qquad\qquad+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\big{)}\end{split}$
Proof.
First, let us prove the estimate for $\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}$. Let $\boldsymbol{x}_{I}\vcentcolon=(y_{{\mathcal{B}}_{h}},u_{I})$ be given by Lemma 4 and 5. Applying the triangle inequality, we have
(41) $\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}\leq\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}+\left\lVert{\boldsymbol{x}_{I}-\boldsymbol{x}_{h}}\right\rVert_{X}\,,$
then the term $\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}$ can be estimated applying (36) and (34), i.e.
(42) $\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}&=\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}+\left\lVert{u-u_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\\ &\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\end{split}$
where $\xi\lesssim\zeta$ denotes that there exists a constant $C$ independent of $h$ such that $\xi\leq C\zeta$. This notation is widely used in the following. On the other hand, let $\boldsymbol{\varepsilon}_{h}=\boldsymbol{x}_{I}-\boldsymbol{x}_{h}$. Noticing that $\boldsymbol{\varepsilon}_{h}\in X_{h,0}$ thanks to Lemma 5, we can apply the coercivity of ${\mathcal{A}}_{h}$ on $X_{h,0}$ (26) and, adding and subtracting terms, we obtain
(43) $\alpha_{\ast}\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}^{2}$ $\leq{\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$
$={\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)+{\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)\mp{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)\,.$
Exploiting the continuity of ${\mathcal{A}}_{h}$ and (42), the first term of (43) can be estimated as
(44) ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$
Then, the term ${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)$ of (43) can be bounded exploiting the definition of $\Pi^{0,E}_{k}$, the Cauchy-Schwarz inequality and (31), i.e.
${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)$ $={m}_{h}\left(y,y_{{\mathcal{B}}_{h}}-y_{h}\right)-\left(y,y_{{\mathcal{B}}_{h}}-y_{h}\right)_{\Omega_{\text{obs}}}$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\qquad\qquad\qquad\qquad-\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y,(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}y-y,(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\lesssim h^{s+1}\left\lvert{y}\right\rvert_{s+1,\Omega}h\left\lVert{\nabla(y_{{\mathcal{B}}_{h}}-y_{h})}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}$
$\lesssim h^{s+2}\left\lvert{y}\right\rvert_{s+1,\Omega}\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$
The remaining term in (43), ${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$, can be rewritten as
${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$ $={\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{B}}\left(p,\boldsymbol{\varepsilon}_{h}\right)$
$\quad\quad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$={\mathcal{B}}_{h}\left(p_{I},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{B}}\left(p,\boldsymbol{\varepsilon}_{h}\right)$
$\qquad\qquad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$={a}_{h}\left(y_{{\mathcal{B}}_{h}}-y_{h},p_{I}\right)-{a}\left(y_{{\mathcal{B}}_{h}}-y_{h},p\right)$
$\qquad\qquad+\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}$
$\qquad\qquad\qquad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
where we apply Problems 3 and 4, Lemma 3 to define $p_{I}\in Y_{h}$ interpolating $p$, Lemma 5 and the definition of $\Pi^{0,E}_{k}$. Then, the term ${a}_{h}\left(y_{{\mathcal{B}}_{h}}-y_{h},p_{I}\right)-{a}\left(y_{{\mathcal{B}}_{h}}-y_{h},p\right)$ is estimated as
(45) ${a}_{h}\left(y_{{\mathcal{B}}_{h}}-y_{h},p_{I}\right)-{a}\left(y_{{\mathcal{B}}_{h}}-y_{h},p\right)\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert y_{{\mathcal{B}}_{h}}-y_{h}\right\rvert\right\rvert\right\rvert_{Y}\,,$
following the same steps applied in the Proof of Lemma 5 (see (39)). The term $\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}$ is estimated using the Cauchy-Schwarz inequality and (33), obtaining
(46) $\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}\lesssim h^{s+1}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lVert{u_{I}-u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\,.$
The last term of (4) is estimated using the Cauchy-Schwarz inequality and (31) as
(47) $\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\qquad\qquad=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\qquad\qquad\lesssim h^{\tau}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}h\left\lVert{\nabla(y_{{\mathcal{B}}_{h}}-y_{h})}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}$
$\qquad\qquad\lesssim h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\left\lvert\left\lvert\left\lvert y_{{\mathcal{B}}_{h}}-y_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$
Hence, coming back to (4), we obtain that
(48) ${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)\lesssim\left(h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\right)\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$
Substituting in (43) the estimates (44), (4) and (48), we get
(49) $\begin{split}\left\lVert{\boldsymbol{x}_{I}-\boldsymbol{x}_{h}}\right\rVert_{X}&\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\\ &\qquad\qquad\qquad+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,.\end{split}$
Hence, considering also (41) and (42), we have the estimate for $\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}$, i.e.
(50) $\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}&\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\\ &\qquad\qquad\qquad+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,.\end{split}$
Now, let us focus on the estimate of $\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}$. Let $p_{I}\in Y_{h}$, interpolating $p$, be given by Lemma 3. Applying the triangle inequality, we have
(51) $\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\left\lvert\left\lvert\left\lvert p-p_{I}\right\rvert\right\rvert\right\rvert_{Y}+\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}$
From (33), we obtain that
(52) $\left\lvert\left\lvert\left\lvert p-p_{I}\right\rvert\right\rvert\right\rvert_{Y}\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\,.$
On the other hand, starting from the inf-sup condition (27), we have
(53) $\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\frac{1}{\beta_{\ast}}\sup_{\boldsymbol{\varsigma}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)}{\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}}\,.$
Let us estimate ${\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)$ for any given $\boldsymbol{\varsigma}_{h}\in X_{h}$. First, adding and subtracting ${\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)$, defining $\boldsymbol{\varsigma}_{h}=(w_{h},v_{h})$ we get
(54) ${\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)$ $={\mathcal{B}}_{h}\left(p_{I},\boldsymbol{\varsigma}_{h}\right)\mp{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)$
$={a}_{h}\left(w_{h},p_{I}\right)-{a}\left(w_{h},p\right)+\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}$
$\qquad\qquad\qquad\qquad\qquad+{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)\,.$
The first two terms are estimated as
(55) ${a}_{h}\left(w_{h},p_{I}\right)-{a}\left(w_{h},p\right)\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$
following, as previously, the same steps applied in the Proof of Lemma 5 (see (39)). Then, applying the Cauchy-Schwarz inequality, a trace inequality together with a Poincarè inequality and (33), the term $\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}$ in (54) is bounded as
(56) $\begin{split}\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}&\leq\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{p-p_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\\ &\lesssim\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{\nabla(p-p_{I})}\right\rVert_{\mathrm{L}^{2}\!(E)}\\ &\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\,.\end{split}$
The last two terms of (54), applying Problems (3) and (4), become
(57) $\begin{split}{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)&=-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)+\left(y_{\mathrm{d}},w_{h}\right)_{\Omega_{\text{obs}}}+{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)\\ &\qquad\qquad\qquad\qquad-\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},\Pi^{0,E}_{k}w_{h}\right)_{E}\,,\end{split}$
where applying the Cauchy-Schwarz inequality and (31), we get
(58) $\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},w_{h}-\Pi^{0,E}_{k}w_{h}\right)_{E}$
$\qquad\qquad\qquad=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},w_{h}-\Pi^{0,E}_{k}w_{h}\right)_{E}$
$\qquad\qquad\qquad\leq\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{w_{h}-\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
$\qquad\qquad\qquad\lesssim h^{\tau}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}h\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\lesssim h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$
On the other hand, the term ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ in (57) is estimated adding and subtracting ${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$, applying the definition of ${\mathcal{A}}_{h}$ and ${\mathcal{A}}$, the definition and the continuity of $\Pi^{0,E}_{k}$ together with the Cauchy-Schwarz inequality, a Poincarè inequality and (31), and (50), i.e.
(59) ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$ $={\mathcal{A}}_{h}\left(\boldsymbol{x}_{h}-\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)+{\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}(y_{h}-y),\Pi^{0,E}_{k}w_{h}\right)_{E}$
$\quad+\alpha\left(u_{h}-u,v_{h}\right)_{\Gamma_{C}}$
$\quad\quad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}w_{h}\right)_{E}-\left(y,w_{h}\right)_{E}$
$\leq\left\lVert{y-y_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
$\quad+\alpha\left\lVert{u-u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}$
$\quad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y-\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}w_{h}-w_{h}\right)_{E}$
$\lesssim\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}+h^{s+1}\left\lvert{p}\right\rvert_{s+1}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$
$\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\right)\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}\,$
$\quad\quad+\left(h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\right)\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}\,.$
Finally, considering together (55), (56), (58) and (59), and substituting in (53), we obtain the estimate
$\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,,$
that together with (50) gives the thesis (40). ∎
5. Numerical Results
In this section, we present two numerical tests, that are solved with the proposed scheme using the dofi-dofi stabilization defined in Remark 2. The first test assess the numerical rate of convergence, while the second one explores the choice of the stability parameter $\sigma$. In the framework of the second test, we propose an alternative scheme designed to avoid issues related to the choice of $\sigma$.
5.1. Test 1
The convergence test is carried out on a $\Omega=(0,1)^{2}$. The boundary is partitioned in $\Gamma_{N}=\emptyset$, $\Gamma_{C}=\{0\}\times(0,1)$ and $\Gamma_{D}=\partial\Omega\setminus\Gamma_{C}$. The spatial coordinates are denoted with $x=(x_{1},x_{2})$. Figure 1 depicts a schematic representation of the domain. In this case, we deal with distributed observation, i.e. $\Omega_{\text{obs}}=\Omega$.
$(0,0)$$(1,0)$$(1,1)$$(0,1)$$\Omega$$\Gamma_{D}$$\Gamma_{C}$
Figure 1. Test 1: schematic representation of the domain.
Let
$y(x_{1},x_{2})=(1-x_{1})\sin(\pi x_{2}),\quad u(x_{1},x_{2})=\sin(\pi x_{2})$
and
$p(x_{1},x_{2})=\sin(\pi x_{2})\cos\left(\frac{\pi}{2}x_{1}\right)$
be the exact solutions of Problem 3 with $\alpha=1$, $k(x)\equiv\gamma(x)\equiv 1$ and
$f(x_{1},x_{2})$ $=(1-x_{1})\sin(\pi x_{2})(1+\pi^{2})\,,$
$y_{\mathrm{d}}(x_{1},x_{2})$ $=\sin(\pi x_{2})\left[(1-x_{1})+\left(\frac{5\pi^{2}}{4}+1\right)\cos\left(\frac{\pi}{2}x_{1}\right)\right].$
(a) Cartesian.
(b) Star.
(c) Polymesher.
Figure 2. Test 1: meshes.
(a) Cartesian, $\mathrm{L}^{2}$ error.
(b) Star, $\mathrm{L}^{2}$ error.
(c) Polymesher, $\mathrm{L}^{2}$ error.
(d) Cartesian, e-error.
(e) Star, e-error.
(f) Polymesher, e-error.
Figure 3. Test 1: convergence plots for the state variable $y$.
(a) Cartesian, $\mathrm{L}^{2}$ error.
(b) Star, $\mathrm{L}^{2}$ error.
(c) Polymesher, $\mathrm{L}^{2}$ error.
(d) Cartesian, e-error.
(e) Star, e-error.
(f) Polymesher, e-error.
Figure 4. Test 1: convergence plots for the adjoint variable $p$.
(a) Cartesian, $\mathrm{L}^{2}$ error.
(b) Star, $\mathrm{L}^{2}$ error.
(c) Polymesher, $\mathrm{L}^{2}$ error.
Figure 5. Test 1: convergence plots for the control variable $u$.
We solve Problem 4 with the stability parameter $\sigma=1$ for $k=1,\ldots,4$ on three different grids, depicted in Figure 2, obtaining the approximated solutions $y_{h},u_{h}$ and $p_{h}$. The considered meshes are: a tessellation of squares (named Cartesian, see Figure 2(a)), a mesh of non-convex octagons and nonagons (named Star, Figure 2(b)) and mesh of convex polygons obtained exploiting Polymesher [35] (named Polymesher, Figure 2(c)). For each mesh, we consider four refinement levels to show the convergence properties stated in Theorem 1, computing the following errors:
  • the $\text{L}^{2}$ error measures
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\Pi^{0,E}_{k}y_{h}-y}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}},\quad\displaystyle\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\Pi^{0,E}_{k}p_{h}-p}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}}$
    and
    $\left\lVert{u_{h}-u}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})},$
  • the energy error (e-error) measures
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\sqrt{\mathcal{K}}\left(\nabla\Pi^{\nabla,E}_{k}y_{h}-y\right)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\sqrt{\gamma}(\Pi^{0,E}_{k}y_{h}-y)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}},$
    and
    $\sqrt{\sum_{E\in\mathcal{M}_{h}}\left\lVert{\sqrt{\mathcal{K}}\left(\nabla\Pi^{\nabla,E}_{k}p_{h}-p\right)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\sqrt{\gamma}(\Pi^{0,E}_{k}p_{h}-p)}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}}.$
We report the trend of the L2 and the energy error for the state variable $y$ and the adjoint variable $p$ in Figures 3 and 4. The convergence plot for the L2 error of the control variable $u$ is depicted in Figure 5. In the legends, we specify the average convergence rate $m$ with respect to the mesh diameter $h$. We notice that the rates match the expected values stated in Theorem 1, even though some results for order $k=4$ are affected by machine precision issues.
5.2. Test 2 (sensitivity study on the stabilization parameter $\sigma$)
$\Omega$$\Omega_{1}$$\Omega_{2}$$\Gamma_{D}$$\Gamma_{N}$$\Gamma_{C}$$\Gamma_{C}$$(0,0)$$(2,0)$$(2,1)$$(2,0)$
Figure 6. Test 2: schematic representation of the spatial domain.
(a) Polymesher60.
(b) Polymesher120.
Figure 7. Test 2: meshes.
(a) State.
(b) Adjoint.
(c) Control at $y=0$, $x\in[1,2]$.
(d) Control at $y=1$, $x\in[1,2]$.
Figure 8. Test 2: reference solution.
(a) State, $L^{2}$ error.
(b) Adjoint, $L^{2}$ error.
(c) Control, $L^{2}$ error.
(d) State, e-error.
(e) Adjoint, e-error.
Figure 9. Polymesher60. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
(a) State, $L^{2}$ error.
(b) Adjoint, $L^{2}$ error.
(c) Control, $L^{2}$ error.
(d) State, e-error.
(e) Adjoint, e-error.
Figure 10. Polymesher120. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
We propose a numerical study of VEM for several stabilization parameters $\sigma$, see Remark 2, for the following OCP on a setting inspired by [29]. The physical domain is $\Omega=(0,2)\times(0,1)$. The desired state is observed in $\Omega_{\text{obs}}=\Omega_{1}\cup\Omega_{2}$, for $\Omega_{1}=(1,2)\times(0.8,1)$ and $\Omega_{2}=(1,2)\times(0,0.2)$. The control boundary $\Gamma_{C}$ is $((1,2)\times\{1\})\cup((1,2)\times\{0\}).$ Homogeneous Neumann boundary conditions are imposed over $\{2\}\times(0,2)$. On $\Gamma_{D}=\partial\Omega\setminus(\Gamma_{C}\cup\Gamma_{N})$, we impose non-homogeneous Dirichlet boundary condition $g=1$. We remark that, even if the state equation features non-homogeneous Dirichlet boundary conditions, the adjoint equation still features homogeneous Dirichlet conditions, respectively [22]. The domain is represented in Figure 6. We solve Problem 4 with $\alpha=0.07$, $k(x)\equiv 1/12$, $\gamma(x)=1$, $f=0$ and $y_{\mathrm{d}}\equiv 2.5$ on two meshes, depicted in Figure 7: (i) Polymesher60, a conforming mesh with $60$ polygons built using Polymesher [35] and (ii) Polymesher120, another conforming Polymesher mesh of $120$ polygons. We solve the problem for $k=1,2,3,4$ and $\sigma=10^{-3},10^{-2},10^{-1},1,10,100,1000$. To test the accuracy of the VEM solution, we compare it to a reference finite element solution over a triangular mesh of $55000$ elements. The reference solution has been simulated though FEniCS library [26] and is depicted in Figure 8. From Figures 9 and 10, we can observe a variability of the errors both in term of L2 and energy norm. Let us focus on Figure 9, referring to the results obtained on Polymesher60. The minimum of the error in the L2 norm (top row) shifts from left to right increasing the polynomial degree. For example, if we consider state and adjoint and $k=1,2$ the optimal $\sigma$ is $10^{-1}$, while for $k=3,4$ the optimal value is $1$. For the control we see that $\sigma=1$ is optimal for $k=1,3,4$, while for $k=2$ the minimum L2 error is reached for $\sigma=10^{-1}$. From this, it is clear that different variables may give optimal error values for different $\sigma$. This is even more evident for the energy norm (bottom row) for state and adjoint for $k=2,3,4$: for the state the optimal value is $\sigma=1$, while for the adjoint variable smaller values of $\sigma$ would be beneficial. For $k=1$, the state optimal value in energy norm is reached for $\sigma=10^{-1}$, while for the adjoint is related to $\sigma=10^{3}$. Similar conclusions can be drawn from Figure 10, related to the results obtained on Polymesher120. We highlight that the optimal stabilization parameter may also vary in terms of the mesh we are using. For example, if we compare Figures 9(e) and 10(e) for $k=1$, the adjoint optimal error is reached for $\sigma=10^{3}$ and $\sigma=10^{-2}$, respectively.
5.3. Test 2 (structure preserving VEM)
In this section we tackle the numerical setting discussed in Section 5.2 through the employment of structure preserving VEM [9], to offer an alternative scheme that does not rely on an arbitrary non-polynomial stabilizing bilinear form. At the time of writing this paper, the scheme in [9] is available only for $k=1$. Under assumptions A.1 and A.2, for each $E\in\mathcal{M}_{h}$, let ${a}^{E}_{h}:\mathrm{H}^{1}\!(E)\times\mathrm{H}^{1}\!(E)\to\mathbb{R}$ be the local bilinear form defined as, for all $w_{h},q_{h}$ in $\mathrm{H}^{1}\!(E)$
(60) ${\tilde{a}}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\mathcal{K}\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h},\hat{\Pi}^{0,E}_{\ell_{E}}\nabla q_{h}\right)_{E}+\left(\gamma\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h},\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}q_{h}\right)_{E},$
where, given $\ell_{E}\in\mathbb{N}$, $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla:\mathrm{H}^{1}\!(E)\rightarrow\boldsymbol{\mathrm{curl}}\,\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$ is the $\mathrm{L}^{2}$-projection operator of functions in $\mathrm{H}^{1}\!(E)$, defined, for all $u\in\mathrm{H}^{1}\!(E)$, by the orthogonality condition
(61) $\left(\hat{\Pi}^{0,E}_{\ell_{E}}\nabla u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}=\left(\nabla u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}\quad\forall p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)\,,$
with $\boldsymbol{\mathrm{curl}}\,p=\left(\frac{\partial p}{\partial y},-\frac{\partial p}{\partial x}\right)$ for any $p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$, and $\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}:\mathrm{H}^{1}\!(E)\to\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$ is the operator defined, for all $u\in\mathrm{H}^{1}\!(E)$, by
(62) $\left(\boldsymbol{\mathrm{curl}}\,\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}$ $=\left(\nabla u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}\quad\forall p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)\,,$
$\int_{E}\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u-u$ $=0\,.$
We remark that considering together the two projectors the following relation holds true
(63) $\left(\hat{\Pi}^{0,E}_{\ell_{E}}\nabla u-\boldsymbol{\mathrm{curl}}\,\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u,\boldsymbol{\mathrm{curl}}\,p\right)_{E}=0\quad\forall p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)\,,$
hence $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla u=\boldsymbol{\mathrm{curl}}\,\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u$ for any $u\in\mathrm{H}^{1}\!(E)$. On $\ell_{E}$ we assume:
  1. H.1
    $\ell_{E}$ is chosen such that there exists a set of degrees of freedom for $p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$ containing the integral mean $\frac{1}{\left\lvert{e}\right\rvert}\int_{e}p$ on each edge $e\in{\mathcal{E}_{E}}$.
We refer the reader to [9, Theorem 3] for conditions over $E$ to guarantee H.1. We here propose an extension of [9, Theorem 2].
Theorem 2.
Assuming A.1, A.2 and H.1, for every $E\in\mathcal{M}_{h}$, the discrete bilinear form (60) is coercive on $Y_{h}(E)$, i.e. there exists $\eta^{E}_{*}>0$, independent of $h_{E}$, such that
(64) ${\tilde{a}}^{E}_{h}\left(w_{h},w_{h}\right)\geq\eta^{E}_{*}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}\quad\forall w_{h}\in Y_{h}(E)\,.$
Proof.
Let $w_{h}\in Y_{h}(E)$ be given. Theorem 2 in [9] states that, under the current assumptions,
(65) $\left(\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h},\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h}\right)_{E}\gtrsim\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,.$
Now, we want to show that $\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}u}\right\rVert_{\mathrm{L}^{2}\!(E)}\gtrsim\left\lVert{u}\right\rVert_{\mathrm{L}^{2}\!(E)}$. Considering the definition of the projectors $\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}$ and $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla$ and (63), $\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}=\left\lVert{\hat{\Pi}^{0,E}_{\ell_{E}}\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$ holds true. Hence, we can infer that $\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\gtrsim\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$.
Let us recall that
$\Pi^{0,E}_{0}w_{h}=\left\lvert{E}\right\rvert^{-1}\int_{E}w_{h}=\left\lvert{E}\right\rvert^{-1}\int_{E}\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}\,,$
where the last equality is due to (62). It implies that
$\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$ $=\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2},$
$\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$ $=\left\lVert{w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,.$
Moreover, notice that the following Poincaré inequality
$\left\lVert{w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\lesssim h_{E}\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
and the inverse inequality
$\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\gtrsim h_{E}\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
hold true. Hence, we obtain
(66) $\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$ $=\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$
$\gtrsim h^{2}_{E}\left\lVert{\nabla\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$
$\gtrsim h_{E}^{2}\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}$
$\gtrsim\left\lVert{w_{h}-\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\Pi^{0,E}_{0}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\gtrsim\left\lVert{w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\,.$
Hence, $\left\lVert{\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$ is a norm for each $w_{h}\in Y_{h}(E)$. Finally, under the hypothesis on $\mathcal{K}$ and $\gamma$, it is easy to see that (65) and (66) imply that the thesis holds true. ∎
(a) State, $L^{2}$ error
(b) Adjoint, $L^{2}$ error
(c) Control, $L^{2}$ error
(d) State, e-error
(e) Adjoint, e-error
Figure 11. Polymesher60. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
(a) State, $L^{2}$ error.
(b) Adjoint, $L^{2}$ error.
(c) Control, $L^{2}$ error.
(d) State, e-error.
(e) Adjoint, e-error.
Figure 12. Polymesher120. Errors with respect to reference solutions, for $\sigma=10^{-3},10^{-2},10^{1},1,10,100,1000$ and $k=1,2,3,4$.
By Theorem 2, we can solve Problem 4 substituting for each $E\in\mathcal{M}_{h}$ ${a}^{E}_{h}$ with ${\tilde{a}}^{E}_{h}$ and retrieve the well-posedness of the method (as in Proposition 1) and the error estimates (as in Theorem 1). In practice, for the computation of ${\tilde{a}}^{E}_{h}$ we exploit the QR-based algorithm presented in [9], that returns the matrix $\boldsymbol{\Pi}=\left\{\pi_{i,j}\right\}_{ij}$ such that
(67) $\hat{\Pi}^{0,E}_{\ell_{E}}\nabla\phi_{j}=\sum_{i=1}^{\dim\mathbb{P}_{\ell_{E}+1}\!\left(E\right)-1}\pi_{ij}\,\boldsymbol{\mathrm{curl}}\,m_{i}\,,\quad\forall j=1,\ldots,N_{E}$
where $\{\phi_{j}\}_{j=1}^{N_{E}}$ is a set of basis functions of $Y_{h}(E)$ and $\{m_{i}\}_{i=1}^{\dim\mathbb{P}_{\ell_{E}+1}\!\left(E\right)}$ is a set of basis functions of $\mathbb{P}^{0}_{\ell_{E}+1}\!\left(E\right)\vcentcolon=\left\{p\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right):\int_{E}p=0\right\}$. Then, using (63) we get
(68) $\Pi^{\boldsymbol{\mathrm{curl}}\,,E}_{\ell_{E}+1}\phi_{j}=\left\lvert{E}\right\rvert^{-1}\int_{E}\phi_{j}+\sum_{i}\pi_{ij}\,m_{i}\,.$
Hence, the computation of the mass term in ${\tilde{a}}^{E}_{h}$ does not require an additional computational cost.
Now, we present the results obtained with this method. We compare the errors with the ones discussed in Section 5.2 for $k=1$. From Figures 11 and 12 we see that the errors made by structure-preserving VEM (named stab-free in legends) are quite close to the optimal errors made by stabilized VEM, independently of the solution and of the mesh. This suggests that this method is a valid alternative to avoid issues related to the choice of the stability parameter.
6. Conclusions
In this work, we focus on the VEM approximation of Neumann boundary OCPs. We address the problem in saddle point formulation, providing a rigorous a priori error estimation for all the variables at hand. To the best of our knowledge, this is the first time that the analysis is proposed in the saddle point framework for arbitrary polynomial degrees. A numerical test based on analytic solutions confirmed the expected theoretical convergence rates. A second test is proposed: we show the accuracy of the method over a ground truth solution based on FEM discretization. For this latter experiment, we analyze the role of the VEM stabilization and its influence on the approximation error of each component for this coupled problem. We observe mesh-dependency effects and variabilities between the three variables of the system. To circumvent this issue we also propose a structure preserving strategy for polynomial degree $k=1$.
This contribution is meant to be a bridge between the theoretical and applied aspects of the problem, where no exact solution is known. Indeed, this work is a first step towards more complex applications. Indeed, there are many aspects to be investigated to better show the VEM potential for such coupled problems, for example more complicated geometrical structures and models. Moreover, a further interesting potential development would be the extension of theoretical and numerical analysis to other type of OCPs, with different control boundary conditions and constraints.
Acknowledgements
The authors acknowledge the fundings from INdAM - GNCS Projects (CUP: E53C23001670001). AB and FM acknowledge funding by the Italian Ministry of University and Research (MUR) through the project MUR-M4C2-1.1-PRIN 2022 CUP: E53D23005820006. AB ackowledges funding by the European Union through Next Generation EU, M4C2, PRIN 2022 PNRR project CUP: E53D23017950001 and PNRR M4C2 project of CN00000013 National Centre for HPC, Big Data and Quantum Computing (HPC) CUP: E13C22000990001. MS thanks the “20227K44ME - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning (FaReX)” project – funded by European Union – Next Generation EU within the PRIN 2022 program (D.D. 104 - 02/02/2022 Ministero dell’Università e della Ricerca). This manuscript reflects only the authors’ views and opinions and the Ministry cannot be considered responsible for them.